OpenStudy (anonymous):

I kinda need help on this one..... =The system of equations is coincident. What are the missing values? (

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

what is your meaning of the word coincident?

OpenStudy (anonymous):

@suckerofmath

OpenStudy (danjs):

to me it means , that you have 2 equations that describe the same line

OpenStudy (danjs):

your equations, did not showup in the question,

OpenStudy (anonymous):

The lines are overlapped. Plus hears the equation. 4x+5y=8 8x+_y=_

OpenStudy (anonymous):

Heres*

OpenStudy (danjs):

So both equations are the same line , it looks like the second equation is a constant of 2 x the first equations

OpenStudy (danjs):

you can multiply a line equation through by a constant number, and it doesnt change the line at all

OpenStudy (danjs):

4x + 5y = 8 is the same line as 8x + 10y = 16

OpenStudy (danjs):

Everything is multiplied by 2, the equation of the line is the same as the first

OpenStudy (danjs):

12x + 15y = 24 (3x) 16x + 20y = 32 (4x) . . . all the same lines

OpenStudy (anonymous):

Oh okay